• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Z.-E. Meziani

    Log entry time 16:02:05 on April20,2004

    Entry number 118662

    keyword=Shift Summary

    Day Shift
    Began the shift waiting for the beam back.
    8:24 beam back
    8:28 start Production run # 2578 741 K
    9:00 MCC takes the beam back for beam tests to 12:00pm
    9:11 Moved target to empty position requested by MCC
    9:50 MCC requests 12C in position for their tests
    11:15 MCC sending beam at 8 uA and masking the ion chambers as pasrt of their tests
    12:01 Asked MCC for beam as planned.
    Their estimate they still need an hour to troubleshoot.
    13:15 MCC called. They will be sending beam back at
    the highest possible current to check whether their fixes
    13:37 Beam in the Hall but still in testing mode.
    14:32 MCC is still in testing mode.
    16:00 Shift End. MCC needs another hour.